Acceleration is sometimes expressed in multiples of g's where g=9.8 m/s2 . At take off, a jet fighter plane can start from rest and accelerate down the runway to cover .5 mile in 13.5s. on average, how many g's does the pilor experience?
Answer #1

Final Velocity , s = ut+1/2**t^2,

s = 0.5 miles = 804.672 m

804.672 = 0.5*a*13.5)^2

a = 8.83 m/s^2 = 0.901 g